# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v4

@@ 1 SERVER
2       for $t\in[0,1,\cdots,T-1]$  do
3             for $i\in[1,2,\cdots,N]$ parallel  do
@@ 7             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
12       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
@@ 16             Compute $C^{i,j}_{t}$ by Equation (13);
17             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
@@
19      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
@@ 25            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
V-C Overall Algorithm
@@
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in Section V-B, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
VI Finding Optimal Fairness Parameter
@@
VI-A Convergence Analysis
Assumption 1
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3
Theorem 1
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (16). For the detailed proof, please refer to Appendix A-A.
VI-B Optimal Fairness Parameter $\lambda^{*}$
Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (16) to obtain a function in terms of $\lambda$,
@@ VI-B Optimal Fairness Parameter $\lambda^{*}$
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
@@
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l=1,2,3$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
VII Privacy Analysis
Since each client incurs the same privacy loss in every round, we evaluate privacy loss from the perspective of an individual client. Accordingly, we analyze the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 2
@@ VII Privacy Analysis
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4.
VII-A Proof of Theorem 2
Lemma 1
(Conversion from RDP to DP [3]). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
As Algorithm 2 shown, the $i$-th client uploads model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. Both $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the private training data of the client, and thus both require differential privacy protection, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
VII-A1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 3
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 3
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (25)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (26)
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work [37] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (27)
Definition 4
(Composition of RDP[36]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 3 and Definition 4, the Theorem 3 is obtained.
VII-A2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
Theorem 4
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
VII-A3 Proof Loss of FedFDP.
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 2 is hold.
VIII Experiments
TABLE II: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
@@ VIII Experiments
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
TABLE III: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.

### v6

@@ 1 SERVER
2 for $t=0,1,\cdots,T-1$ do
3    for $i=1,2,\cdots,N$ parallel do
@@ 7    $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
12 $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
@@ 16      Compute $C^{i,j}_{t}$ by Equation (13);
17    $\mathbf{\tilde{g}}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
@@
19 $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\mathbf{\tilde{g}}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
@@ 25    $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
Figure 1: FedFDP framework. Besides the privately computed local model $\mathbf{w}_{t+1}^{i}$ from the “Fair-Clipping”, client $i$ is also required to upload the private loss $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
4.3 Overall Algorithm
@@
  1. 1.
    DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\mathbf{\tilde{g}}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\mathbf{\tilde{g}}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  2. 2.
    DP for local loss (lines 2-2): As described in Section 4.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
5 Finding Optimal Fairness Parameter
@@
All assumptions and detailed procedures for the convergence analysis are provided in Appendix 0.A.2, leading to the following result:
Theorem 5.1(Convergence rate of FedFDP   (Simplified version of Theorem 0.A.1))
$\mathbb{E}[F(\mathbf{w}_{T})]-F^{*}$ $\leq\mathcal{O}\left(\frac{LG^{2}\lambda^{2}}{\mu^{2}T}\right)+\mathcal{O}\left(\frac{L^{2}\Gamma\lambda}{\mu^{2}T}\right)$
$\quad+\mathcal{O}\left(\frac{L\sigma^{2}C^{2}d}{\mu^{2}T\lambda}\right)+\mathcal{O}\left(\frac{L\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}}{T}\right)\,$, (16)
where $L$, $\mu$, and $G$ are parameters defined in Assumptions 13, and $\Gamma$ is a measure of heterogeneity; see Appendix 0.A.1 for details.
As demonstrated in (5.1), the fairness parameter $\lambda$ ($\lambda>0$) exhibits convex influence on the convergence rate. The analytical solution of this relationship will be obtained via extremum analysis in Section 5.1.
5.1 Optimal Fairness Parameter $\lambda^{*}$
Assuming $\Delta_{i}^{j}$ is bounded as $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (5.1) to obtain a function in terms of $\lambda$:
@@ 5.1 Optimal Fairness Parameter $\lambda^{*}$
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, we have $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
@@
$x_{1}=\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}},$
$x_{2}=\omega\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\omega^{2}\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}},$
$x_{3}=\omega^{2}\sqrt[3]{-\frac{a_{7}}{2}+\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}+\omega\sqrt[3]{-\frac{a_{7}}{2}-\sqrt{\frac{a_{7}^{2}}{4}+\frac{a_{6}^{3}}{27}}}.$
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l\in\{1,2,3\}$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $\mathrm{m}$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{\mathrm{m}}(\lambda_{\mathrm{m}})$.
6 Privacy Analysis
At the beginning of privacy analysis, we clarify that: (1) FedFDP does not assume Secure Aggregation (SecAgg)[4], allowing the server to observe individual client uploads; (2) we guarantee sample-level DP via per-sample clipping to protect individual data records; and (3) SecAgg is orthogonal to our method—while it hides per-client updates, it cannot replace DP in preventing the aggregated model from memorizing sensitive data.
As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of client $i$, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
6.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 6.1
where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
Proof. We will prove Theorem 6.1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 3
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))\,$, (24)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)=\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\,,\\ B_{\alpha}(q,\sigma)=\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\,.\end{array}\right$. (25)
with $\vartheta_{0}=\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}=\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta=(1-q)\vartheta_{0}+q\vartheta_{1}$ .
Further, it holds for $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work [32] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$:
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)\,.$ (26)
Definition 4
(Composition of RDP[33]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as (X,Y)(X,Y)(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 3 and Definition 4, the Theorem 6.1 is obtained.
6.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 6.2
The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
6.3 Privacy Loss of FedFDP
Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP. Lastly, we can get the Privacy loss of FedFDP as follows.
Theorem 6.3
$\begin{split}\begin{aligned} (\epsilon^{i},\delta^{i})=&(R_{model}^{i}(\alpha)+R_{loss}^{i}(\alpha)+\ln((\alpha-1)/\alpha)\\ &-(\ln\delta+\ln\alpha)/(\alpha-1),\delta),\end{aligned}\end{split}$ (28)
where $0<\delta<1$, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 6.1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 6.2.
Lemma 1
[…]
7 Experiments
In this section, we demonstrate the effectiveness of FedFair/FedFDP through answering the following three research questions:
  • RQ1 How effective is FedFair / FedFDP under classification tasks with real-world datasets?
  • RQ2 How robust is FedFair / FedFDP across different experimental settings?
  • RQ3 What is the performance of FedFDP under differential hyper-parameters?
To address the aforementioned three RQs, Section 7.1 introduces the default experimental settings used in this study, including the description of baselines and the configuration of the experimental code. Section 7.2 presents comparative experiments conducted on three datasets under default parameters involving FedFair, FedFDP, and six baseline methods to answer RQ1. Section 7.3 addresses RQ2 through experiments on heterogeneity and scalability. Finally, Section 7.4 performs ablation studies on $\lambda$, $C$, $\sigma$, $\epsilon$, and the target accuracy to answer RQ3.
7.1 Default Settings
Prior to addressing the three research questions, an outline of the default experimental configuration is provided.
Baselines. We compared our algorithms against six baseline methods. For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was applied to each baseline algorithm to enable a fair comparison, which include FedAvg[30], SCAFFOLD[17], FedProx[24], FedDyn[2], ALI-DPFL[28] and q-FFL[25].
Table 1: Comparison of Communication Overhead and Transmission Requirements. $d$ represent the dimension of model.
Algorithm
Comm. Cost
(Per Round)
Extra Transmitted
Variables
Description of Overhead
FedAvg $\mathcal{O}(d)$ None Transmits only model parameters/gradients.
FedProx $\mathcal{O}(d)$ None Adds proximal term locally.
q-FFL $\mathcal{O}(d)$ Scalar Sends scalar loss for re-weighting.
SCAFFOLD $\mathcal{O}(2d)$ Control Variates Transmits both model and control variates.
FedDyn $\mathcal{O}(d)$ None Improve SCAFFOLD without control variables.
ALI-DPFL $\mathcal{O}(d)$ Scalar Server broadcast scalar local steps to clients.
FedFDP $\mathcal{O}(d)$ Scalar Transmits scalar loss for fairness/privacy.
Communication Overhead Analysis. Table 1 presents a detailed comparison of the communication complexity per round for FedFDP and other state-of-the-art baselines. While methods like SCAFFOLD introduce control variates to mitigate client drift, they double the communication payload to $\mathcal{O}(2d)$, which can be prohibitive in bandwidth-constrained wireless networks. In contrast, FedFDP maintains a communication complexity of $\mathcal{O}(d)$, aligning with efficient baselines such as FedAvg and FedDyn . Although FedFDP incorporates a fairness-aware mechanism and differential privacy, the additional transmission overhead is restricted to a negligible scalar value (representing the fairness loss), similar to the approach in q-FFL and ALI-DPFL. This demonstrates that FedFDP achieves robust privacy and fairness without compromising communication efficiency.
Tasks setting. By presenting the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [19], FashionMNIST [40], and CIFAR10 [18], we adopted a widely used heterogeneous settings [27, 21] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [27, 38]. We use a 4-layer CNN architecture [30] which consists of two convolutional layers and two fully connected layers as model architecture.
Implementation environment. We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running on Windows 10.
Hyperparameters. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1.0\times 10^{-5}$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
7.2 Effective in Real-world Datasets
To address RQ1, we conducted experiments on three datasets for both FedFair and FedFDP under the default parameter settings described in Section 7.1. The results, presented in Table 2 and Table 3, report the test accuracy ($\%$) and fairness measure ($\Psi$).
Table 2: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
@@ 7.2 Effective in Real-world Datasets
[…]
Table 3: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
@@
[…]
Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%. A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$. Compared to training methods with a fixed learning rate, learning rate decay accelerate the convergence of gradient-based methods [45].
Table 3 presents the experimental results of the FedFDP algorithm compared to baseline methods. In comparison with Table 2, the introduction of DP leads to a slight decrease in accuracy, approximately in the range of $3\%$ - $10\%$. However, the fairness issue becomes more pronounced, with the corresponding metric increasing significantly in magnitude, indicating that DP exerts a substantial influence on fairness in FL. Under a fixed privacy budget of $\epsilon=3.52$, FedFDP significantly enhances fairness while matching the accuracy of the best-performing baseline, achieving improvements of $30.3\%$, $6.7\%$, and $27.8\%$ across three distinct datasets, respectively.
7.3 Robust in Different Settings
To answer RQ2, as shown in Table 4 and Table 5, we extended the heterogeneity setting from the default $\mathrm{Dir}(0.1)$ to $\mathrm{Dir}(0.5)$ and $\mathrm{Dir}(1)$, and scaled the number of clients from the default 10 to 20 and 50 to examine scalability.
Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.